Sorry,Just one or two more questions.Thanks for the last math problem,I understood it!I'm Still so confused about fractions,though.Help Me!!!!!!!
x/3 - y/2 = -10/3
-4x + 5y = 36
The Answer Choices Are (16/3,-2) (-3,14/3) (-4,4) (4,28/3)

I'M SO CONFUSED!!!!!!!!!1111

Answer by fractalier(6550)

Oh I see. You only gave me one equation before...okay from
x/3 - y/2 = -10/3
-4x + 5y = 36
Let us multiply the top equation by twelve. By doing that, not only we will we clear all of the fractions, but the number of x's will match the number of x's in the second equation, enabling us to eliminate x's by adding equations...here goes...
12[x/3 - y/2 = -10/3]=
4x - 6y = -40
Now add that to the second equation:
4x - 6y = -40
-4x + 5y = 36
and we get
-y = -4
and y = 4
Now we plug that value into either equation to find x:
-4x + 5(4) = 36
-4x + 20 = 36
-4x = 16
x = -4
So (-4, 4) is your answer.